40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14776     Entry time: Fri Jul 19 12:50:10 2019
Author: gautam 
Type: Update 
Category: SUS 
Subject: DC bias actuation options for SOS 

Rana and I talked about some (genius) options for the large range DC bias actuation on the SOS, which do not require us to supply high-voltage to the OSEMs from outside the vacuum.

What we came up with (these are pretty vague ideas at the moment):

  1. Some kind of thermal actuation.
  2. Some kind of electrical actuation where we supply normal (+/- 10 V) from outside the vacuum, and some mechanism inside the chamber integrates (and hence also low-pass filters) the applied voltage to provide a large DC force without injecting a ton of sensor noise.
  3. Use the blue piers as a DC actuator to correct for the pitch imbalance --- Kruthi and Milind are going to do some experiments to investigate this possibility later today.

For the thermal option, I remembered that (exactly a year ago to the day!) when we were doing cavity mode scans, once the heaters were turned on, I needed to apply significant correction to the DC bias voltage to bring the cavity alignment back to normal. The mechanism of this wasn't exactly clear to me - furthermore, we don't have a FLIRcam picture of where the heater radiation patter was centered prior to my re-centering of it on the optic earlier this year, so we don't know what exactly we were heating. Nevertheless, I decided to look at the trend data from that night's work - see Attachment #1. This is a minute trend of some ETMY channels from 0000 UTC on 18 July 2018, for 24 hours. Some remarks:

  1. We did multiple trials that night, both with the elliptical reflector and the cylindrical setup that Annalisa and Terra implemented. I think the most relevant part of this data is starting at 1500 UTC (i.e. ~8am PDT, which is around when we closed shop and went home). So that's when the heaters were turned off, and the subsequent drift of PIT/YAW are, I claim, due to whatever thermal transients were at play.
  2. Just prior to that time, we were running the heater at close to its maximum rated current - so this relaxation is indicative of the range we can get out of this method of actuation.
  3. I had wrongly claimed in my discussion with Rana this morning that the change in alignment was mostly in pitch - in fact, the data suggests the change is almost equal in the two DoFs. Oplev and OSEMs report different changes though, by almost a factor of 2....
  4. The timescale of the relaxation is ~20 minutes - what part(s) of the suspension take this timescale to heat up/cool down? Unlikely to be the wire/any metal parts because the thermal conductivity is high? 
  5. In the optimistic scenario, let's say we get 100 urad of actuation range - over 40m, this corresponds to a beam spot motion of ~8mm, which isn't a whole lot. Since the mechanism of what is causing this misalignment is unclear, we may end up with significantly less actuation range as well.
  6. I will repeat the test (i.e. drive the heater and look for drift in the suspension alignment using OSEMs/Oplev) in the afternoon - now I claim the radation pattern is better centered on the optic so maybe we will have a better understanding of what mechanisms are at play.

Also see this elog by Terra.

Attachment #2 shows the results from today's heating. I did 4 steps, which are obvious in the data - I=0.6A, I=0.76A, I=0.9A, and I=1.05A.


In science, one usually tries to implement some kind of interpretation. so as to translate the natural world into meaning.

Attachment 1: heaterPitch_2018.pdf  1.111 MB  Uploaded Fri Jul 19 13:51:09 2019  | Hide | Hide all
heaterPitch_2018.pdf
Attachment 2: Screenshot_from_2019-07-19_16-39-21.png  70 kB  Uploaded Fri Jul 19 19:26:16 2019  | Hide | Hide all
Screenshot_from_2019-07-19_16-39-21.png
ELOG V3.1.3-